OpenStudy (anonymous):

- -2u^2v^4.uv^-2 OVER (u^-3)^3

OpenStudy (anonymous):

Is it this: \[\frac{-2u^2v^4\times(uv)^{-2}}{(u^{-3})^3}\] ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

And what you need? :-)

OpenStudy (anonymous):

I need to factorr I believe

OpenStudy (anonymous):

sorry I need to simplify

OpenStudy (candy13106):

hey keith can u plz help me on mine

OpenStudy (anonymous):

I keep getting 2u^12v^2 but the two should be negative

OpenStudy (anonymous):

Haha alright! well you need to know two laws: \(x^{-a}=\frac{1}{x^a}\) And \((x^a)^b=x^{ab}\) So we can proceed: \[\frac{-2u^2v^4\times(uv)^{-2}}{(u^{-3})^3}=\frac{-2u^2v^4}{(u^{-9})(uv)^2}=\frac{-(u^9)2u^2v^4}{(uv)^2}=\frac{-2u^{11}v^4}{u^2v^2}=-2u^9v^2\]

OpenStudy (anonymous):

Im sorry but that's very confusing

OpenStudy (anonymous):

Haha its okay my bad. So you start with The original equation. Then the second step I remembered that \((u^{-3})^3=u^{-9}\) and that \((uv)^{-2}=\frac{1}{(uv)^2}\) Okay?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

And then the second step I realized \(\frac{1}{u^{-9}}=u^9\)

OpenStudy (anonymous):

Cool?

OpenStudy (anonymous):

Uhhh okay

OpenStudy (anonymous):

yo I just need help with the last step, cause it goes to - (-2u^12v^2) so wouldnt negaitve and negative make positve 2?

OpenStudy (anonymous):

And then the third step I realized that \((u^9)(u^2)=u^{11}\) and that \((uv)^2=u^2v^2\) And the last step I divided everything out

OpenStudy (anonymous):

achaa alright alright thanks ! makes a bit more sense now

OpenStudy (anonymous):

Lol great im glad!

OpenStudy (anonymous):

thank you :)